Center of mass· X-M mAxA+mBXB - -?? onstants g 9.80 m/s2 During a basketball dribble, the basketball as a downward velocity of 2.40 ms just before hitting the floor and bounces back with an upward velocity of 2.20 m/s. The mass of the basketball is 0.620 kg. The contact time between the basketball and the floor is 0.115 s. Find the average force that the floor exerts on the basketball during the bounce. (Include th basketballs weight) Answer: At 30.9 N C. 18.7 N D. 6.08 N B lauhall of mass m

Please write out the full answer!
Answer #1

Avg force is given by:

F = dP/dt

F = m*dV/dt

Given values are:

m = 0.620 kg

dt = 0.115 sec

dV = Vf - Vi = 2.20 - (-2.40) = 4.6 m/sec

So,

Favg = 0.620*4.6/0.115

Favg = 24.8 N